# HumanVid: Demystifying Training Data for Camera-controllable Human Image Animation

### v1

@@ 3.3 CamAnimate
To validate our dataset’s capability for animating humans with moving cameras, we propose a simple baseline for the camera-controllable human image animation task, named CamAnimate. By leveraging CameraCtrl [23]’s advanced camera pose control and Animate Anyone [29]’s character animation framework, CamAnimate ensures consistent and high-quality human video generation with simultaneous human and camera movements. As shown in Fig. 5, it utilizes plücker embeddings to accurately parameterize camera trajectories and incorporates an additional camera pose encoder to encode camera information for the Denoising UNet via zero-convolution [79], while ReferenceNet and a pose guider maintain appearance consistency and pose controllability. By training our method on general human videos with both camera and human movement, these two types of motion can be decoupled in the network and learned by separate modules in an end-to-end manner. In addition to the moving camera setting, CamAnimate can seamlessly handle traditional human image animation given a static camera parameter.
Implementation Details. We use the checkpoint of Stable Diffusion 1.5 [50] to initialize the Denoising UNet and ReferenceNet, and use the weights of ControlNet [80] on OpenPose [15] to initialize the Pose Guider. The camera encoder weights from CameraCtrl [23] are used to initialize our camera encoder. We mix train horizontal and vertical videos with a resolution of (long side, short side) $=(896,512)$, i.e., for horizontal videos $(w,h)=(896,512)$ and for vertical videos $(w,h)=(512,896)$. Each batch only samples either all horizontal or all vertical videos, and the choice between horizontal and vertical videos is made randomly between batches. The setting of such resolution is according to the GPU memory in our experiments, i.e., maximum GPU memory usage with such pairs of batch size and resolution in both stages. We empirically find that such resolution could achieve a balance of visual quality and computational cost. In the first stage, we train all network parameters using a batch size of 8. In the second stage, we freeze the denoising UNet, reference UNet, and pose guider, and only train the camera encoder and motion module. The motion module in the second stage is initialized with the weights from AnimateDiff [22] V3. The frame rate is set to 24, and the batch size is 1. We use 8 NVIDIA A100 GPUs for all training stages and 1 NVIDIA A100 GPU for testing. The first and second stages are trained for 30,000 and 10,000 iterations, respectively, with a learning rate of 1e-5 and AdamW optimizer [33]. Our camera embedding representation is the Plücker embedding from CameraCtrl, which is computed from the camera’s intrinsic and extrinsic parameters. For real internet data, the intrinsic parameters are set using a heuristic value, while the extrinsic parameters are obtained using SLAM [60] methods. For synthetic data, the intrinsic and extrinsic parameters are directly exported.
@@ 4 Experiments
4.1 Evaluation Benchmark
Due to the lack of a unified benchmark for previous methods, the testing protocols for each method have been varied significantly. However, when the form of the samples differs, the disparity between the reference image and the target image can vary greatly, leading to highly inconsistent results for the same method when inferring different reference and target images. Therefore, as the first large-scale dataset, we provide a unified testing protocol for human video generation. Specifically, we use the middle (12th) frame as the reference image, predict frames in the range [1,72) with a stride of 3, resulting in a sequence of 24 frames. Finally, we evaluate each video under this setting using PSNR [67], SSIM [67], LPIPS [81], FID [24], and FVD [61] metrics.
Table 2: Comparison with SOTA on Tiktok dataset with static camera.
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$ FID $\downarrow$
@@ 4.1 Evaluation Benchmark
Figure 6: Qualitative comparisons with preivous SOTA methods on the test set.
We use the last 40 videos from the TikTok dataset [30] out of a total of 340 videos as the test set for evaluation on static camera. Additionally, we provide 40 videos each in both portrait and landscape orientations as a test set for evaluation on moving camera human video generation, sampled from our collected Internet videos. It is worth noting that SSIM, PSNR, and LPIPS are only reliable under static camera conditions without any human turning. For human videos with camera movements, these reconstruction metrics may not be trustworthy, as video generation inherently involves ambiguity, where we prefer generation metrics like FID and FVD.
@@
Ours 0.672 19.534 0.275 732.7 46.06
@@
Ours 0.678 18.939 0.303 792.2 54.02
@@
4.2 Comparison with the State-of-the-Art
In this section, we compare our baseline model with previous state-of-the-art methods, namely Animate Anyone [29], Magic-animate [71] and Champ [86]. As animate anyone is not open-sourced, we use a third-party implementation [^1] https://github.com/MooreThreads/Moore-AnimateAnyone. We use the official implementations for other two methods. As shown in Tab. 2, although our method is trained on videos with moving cameras, it is still able generate high-quality static camera videos on TikTok dataset and achieves best performance in all metrics due to our precise camera control ability. For human videos with camera movement on our test set, previous methods commonly do not consider the camera condition, so they struggle to produce natural videos with camera movements. Such result could be observed from both reconstruction metrics like SSIM, PSNR, and LPIPS and generation metrics like FID and FVD.
@@ 4.2 Comparison with the State-of-the-Art
User-study. We also conduct a user-study to compare our method with previous methods, as shown in Tab 4. We collect 2 videos from Tiktok test set and 8 videos from our test set, and compare our results with other three methods’ results as a ranking question with 4 options. A total of 20 participants take part in our user study and we assign 3, 2, 1 points for the first, second and third method respectively. The final average score is normalized by total points, so the upper bound of this metric is 0.5. We conclude the average points and top-1 preference of each method in Tab 4, which shows a dominate advantage (0.44 points and 0.73 top-1 preference) over previous methods due to their artifacts in appearance, human pose and camera movements.
@@
Qualitative comparisons. In Fig. 6 , we show the qualitative comparison with previous SOTA methods, where the artifacts are highlighted by red boxes. We show that previous SOTA methods may have different artifacts when applied to our challenging evaluation test sets. For example, animate anyone suffers from inaccurate and low-quality appearances of humans. Champ suffers from missing 3D skeletons due to the difficulty of estimating accurate 3D skeletons, especially in a crowded scene. Magic-animate is not able to correctly generate human face expressions due to the ambiguity of face representation of Densepose. Our method is able to accurately generate facial expressions, body pose and background motions from the camera movement.
@@
Qualitative results on in-the-wild cases. In Fig. 7, we further show our methods ability in generating in-the-wild videos with explicit camera movements. We show that we can achieve high-quality and visually pleasant videos that is close to the actual movies filmed by professionals. We hope that our dataset and the baseline method could serve as a solid base for generating movie-level human videos. Please visit our project page [^2] https://humanvid.github.io/ to view more video results.
@@ 5 Conclusion
In this paper, our study addresses the significant challenges in the field of human image animation by introducing a novel combination of a high-quality real-world video dataset and a meticulously crafted synthetic dataset. Our proposed dataset not only enhances the visual quality and controllability of human animations, but also introduces a new benchmark for camera control in human videos. Without bells and whistles, our proposed simple baseline demonstrate superior performance when it is trained on our combined dataset, particularly in scenarios involving complex human and camera motions. We believe that our contributions will pave the way for more transparent and comprehensive evaluations in this field, fostering further advancements and innovations in video and movie production.
Limitations. The annotation of our Internet data heavily rely on pose estimation [72, 20] and SLAM [66, 60] methods, which could introduce noises into the camera and pose annotations.
@@
Acknowledgment. This project is funded in part by Shanghai AI Laboratory (P23KS00020, 2022ZD0160201), CUHK Interdisciplinary AI Research Institute, and the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK.

### v3

@@ 3.3 CamAnimate
To validate our dataset’s capability for animating humans with moving cameras, we propose a simple baseline for the camera-controllable human image animation task, named CamAnimate. By leveraging CameraCtrl [22]’s advanced camera pose control and Animate Anyone [28]’s character animation framework, CamAnimate ensures consistent and high-quality human video generation with simultaneous human and camera movements. As shown in Fig. 4, it utilizes plücker embeddings to accurately parameterize camera trajectories and incorporates an additional camera pose encoder to encode camera information for the Denoising UNet via zero-convolution [81], while ReferenceNet and a pose guider maintain appearance consistency and pose controllability. By training our method on general human videos with both camera and human movement, these two types of motion can be decoupled in the network and learned by separate modules in an end-to-end manner. In addition to the moving camera setting, CamAnimate can seamlessly generate static-camera human videos. Please refer to Sec. A.1 for more implementation details of CamAnimate.
@@ 4 Experiments
Evaluation Benchmark. Due to the lack of a unified benchmark for previous methods, the testing protocols for each method have been varied significantly. However, when the form of the samples differs, the disparity between the reference image and the target image can vary greatly, leading to highly inconsistent results for the same method when inferring different reference and target images. Therefore, as the first large-scale dataset, we provide a unified testing protocol for human video generation. Specifically, we use the middle frame as the reference image, predict frames in the range [1,72) with a stride of 3, resulting in a sequence of 24 frames. Finally, we evaluate each video under this setting using PSNR [70], SSIM [70], LPIPS [82], FID [23], and FVD [63] metrics.
We use the last 40 videos from the TikTok dataset [29] out of a total of 340 videos as the test set for evaluation on static camera. For UBC-Fashion [80], we use the official split with 500 videos for training and 100 videos for testing. Additionally, we provide 40 videos each in both portrait and landscape orientations as a test set for evaluation on moving camera human video generation, sampled from our collected Internet videos. It is worth noting that SSIM, PSNR, and LPIPS are only reliable under static camera conditions without any human turning. For human videos with camera movements, these reconstruction metrics may not be trustworthy, as video generation inherently involves ambiguity, where we prefer generation metrics like FID and FVD.
Table 3: Comparison with SOTA on TikTok and UBC-Fashion dataset with static camera. $\dagger$ means its official implementation is not trained on UBC-Fashion’s training set.
TikTok Test Set SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$ FID $\downarrow$
@@
[…]
Magic-animate [74$\dagger$ 0.602 6.663 0.552 1583.9 118.76
Animate Anyone [28] 0.914 23.163 0.069 345.4 33.77
Champ [88] 0.922 25.269 0.057 269.4 27.35
Ours 0.929 25.921 0.049 256.6 29.30
Figure 5: Qualitative comparisons with previous SOTA methods on the test set.
@@
Ours (1$\times$ batch size) 0.641 18.008 0.309 960.1 77.73
Ours (4$\times$ batch size) 0.672 19.534 0.275 732.7 46.06
@@
Ours (1$\times$ batch size) 0.675 18.081 0.309 816.5 75.67
Ours (4$\times$ batch size) 0.678 18.939 0.303 792.2 54.02
4.1 Comparison with the State-of-the-Art
In this section, we compare our baseline model with previous state-of-the-art methods, namely Animate Anyone [28], Magic-animate [74] and Champ [88]. As animate anyone is not open-sourced, we use a third-party implementation [^1] https://github.com/MooreThreads/Moore-AnimateAnyone. We use the official implementations for other two methods. As shown in Tab. 3, although our method is trained on videos with moving cameras, it is still able generate high-quality static camera videos on TikTok and UBC-Fashion dataset and achieves best performance in almost all metrics due to our precise camera control ability. For human videos with camera movement on our test set shown in Tab. 4, previous methods commonly do not consider the camera condition, so they struggle to produce natural videos with camera movements. Such result could be observed from both reconstruction metrics like SSIM, PSNR, and LPIPS and generation metrics like FID and FVD. In our training set, we found that the performance could be further improved upon Animate Anyone’s training recipe by increasing the effective batch size from $8\times 8=64$ to $4\times 8\times 8=256$ through gradient accumulation, while maintaining 30,000 iterations in stage 1 training. This enhancement suggests our dataset’s superior scale and diversity.
@@ 4.1 Comparison with the State-of-the-Art
Table 6: Comparison with original Animate Anyone trained without camera condition.
TikTok Test Set SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$ FID $\downarrow$
Animate Anyone [28] 0.658 15.954 0.337 1133.1 53.65
Ours 0.778 18.762 0.247 691.8 41.35
Table 7: Comparison of training strategies on different data parts.
TikTok Test Set
Stage 1 w/
Syn. Data
Stage 2 w/
Syn. Data
[…]
Variant 1 $\times$ $\times$ 0.677 15.957 0.333 1066.9 53.08
Variant 2 $\checkmark$ $\checkmark$ 0.734 17.339 0.287 980.3 56.32
Ours $\times$ $\checkmark$ 0.778 18.762 0.247 691.8 41.35
@@
User-study. We also conduct a user-study to compare our method with previous methods, as shown in Tab. 5. We collect 2 videos from Tiktok test set and 8 videos from our test set, and compare our results with other three methods’ results as a ranking question with 4 options. A total of 20 participants take part in our user study and we assign 3, 2, 1 points for the first, second and third method respectively. The final average score is normalized by total points, so the upper bound of this metric is 0.5. We conclude the average points and top-1 preference of each method in Tab. 5, which shows a dominate advantage (0.44 points and 0.73 top-1 preference) over previous methods due to their artifacts in appearance, human pose and camera movements.
@@
Qualitative comparisons. In Fig. 5 , we show the qualitative comparison with previous SOTA methods, where the artifacts are highlighted by red boxes. We show that previous SOTA methods may have different artifacts when applied to our challenging evaluation test sets. For example, animate anyone suffers from inaccurate and low-quality appearances of humans. Champ suffers from missing 3D skeletons due to the difficulty of estimating accurate 3D skeletons, especially in a crowded scene. Magic-animate is not able to correctly generate human face expressions due to the ambiguity of face representation of DensePose. Our method generates accurate facial expressions, body poses, and background motions that correspond to camera movements.
@@
Qualitative results on in-the-wild cases. In Fig. 6, we further show our methods ability in generating in-the-wild videos with explicit camera movements. We show that we can achieve high-quality and visually pleasant videos that is close to the actual movies filmed by professionals. We hope that our dataset and the baseline method could serve as a solid base for generating movie-level human videos. Additional video results are available on the project page [^2] https://humanvid.github.io/.
4.2 Ablation Study
Contribution Decomposition of Dataset and Method. We construct a variant that training the original Animate Anyone on HumanVid dataset without considering camera condition to verify that existing methods could not be directly applied to our dataset. Since most videos in HumanVid have moving cameras, it creates inherent ambiguity for models attempting to learn both camera movement and human movement simultaneously. Thus, the performance of CamAnimate and the variant trained on the same training set shows significant difference, as shown in Tab. 7. It indicates that the camera encoder in our CamAnimate is necessary for training human video generation models on HumanVid.
Training Strategy. Due to limited appearance diversity, synthetic data alone proves insufficient for training effective human image animation models, resulting in non-semantic, uniform textures that lack meaningful structure. While training with Internet videos enables appearance transfer and pose control capabilities, camera control remains challenging. This limitation primarily stems from the difficulty in extracting accurate camera trajectories from Internet videos using COLMAP [54] or SLAM methods [62], as moving subjects often obscure static backgrounds. To address this, we leverage our synthetic data to enhance camera control accuracy and overall visual quality through a two-stage training strategy. In stage 1, we finetune the ReferenceNet, Pose Guider, DenoisingNet, and Camera Encoder. In stage 2, we focus solely on finetuning the Camera Encoder and Temporal Modules in DenoisingNet. This approach effectively combines high-quality Internet videos for human appearance modeling with synthetic videos for improved camera control. As shown in Tab. 7, our final model outperforms both Variant 1 (using only Internet videos) and Variant 2 (using both video types for both training stages).
Camera Trajectory Evaluation. Due to the inaccurate camera trajectories estimated from our generated videos, it is difficult to obtain accurate camera-related metrics (Translation Error and Rotation Error) used in CameraCtrl [22]. The primary reason is that the generated videos in CameraCtrl only have static scenes and moving cameras, where COLMAP [54] is proven to be effective in extracting accurate camera trajectories. However, because we have moving people and moving cameras, such Structure-from-Motion methods will fail in our setting. Thus, we leave it to our future work to quantitatively verify the camera control ability.
@@ 5 Conclusion
In this paper, our study addresses the significant challenges in the field of human image animation by introducing a novel combination of a high-quality real-world video dataset and a meticulously crafted synthetic dataset. Our proposed dataset not only enhances the visual quality and controllability of human animations, but also introduces a new benchmark for camera control in human videos. Without bells and whistles, our proposed simple baseline demonstrate superior performance when it is trained on our combined dataset, particularly in scenarios involving complex human and camera motions. We believe our dataset establishes a foundation for transparent and comprehensive evaluations in this field, facilitating future advances in video and movie production.
Limitations. The annotation of our Internet data heavily rely on pose estimation [75, 19] and SLAM [69, 62] methods, which could introduce noises into the camera and pose annotations. The number of asset and background scenes in the synthetic part is limited when we compare it to real videos. The rendering quality is also worse than the real videos captured by professional cameras.
@@
Acknowledgment. This project is funded in part by Shanghai Artificial Intelligence Laboratory, CUHK Interdisciplinary AI Research Institute, and the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK. This project is supported by RGC Early Career Scheme (ECS) No. 24209224 and CUHK Direct Grants (RCFUS) No. 4055189. We would like to thank Xuekun Jiang for helpful discussion.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 1.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 5.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 3.3 and supplementary material for details.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See supplementary material.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Due to extensive computation cost, we do not report error bars for PSNR, SSIM, LPIPS, FID and FVD.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See supplementary material.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Section 4
    2. (b)
      Did you mention the license of the assets? [Yes] See supplementary material.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See supplementary material.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See supplementary material.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See supplementary material.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Appendix
A.1 Implementation Details of CamAnimate
We use the checkpoint of Stable Diffusion 1.5 [51] to initialize the Denoising UNet and ReferenceNet, and use the weights of ControlNet [81] on OpenPose [13] to initialize the Pose Guider. The camera encoder weights from CameraCtrl [22] are used to initialize our camera encoder. We mix train horizontal and vertical videos with a resolution of (long side, short side) $=(896,512)$, i.e., for horizontal videos $(w,h)=(896,512)$ and for vertical videos $(w,h)=(512,896)$. Each batch only samples either all horizontal or all vertical videos, and the choice between horizontal and vertical videos is made randomly between batches. The setting of such resolution is according to the GPU memory in our experiments, i.e., maximum GPU memory usage with such pairs of batch size and resolution in both stages. We empirically find that such resolution could achieve a balance of visual quality and computational cost. In the first stage, we train all network parameters using a batch size of 8. In the second stage, we freeze the denoising UNet, reference UNet, and pose guider, and only train the camera encoder and motion module. The motion module in the second stage is initialized with the weights from AnimateDiff [21] V3. The frame rate is set to 24, and the batch size is 1. We use 8 NVIDIA A100 GPUs for all training stages and 1 NVIDIA A100 GPU for testing. The first and second stages are trained for 30,000 and 10,000 iterations, respectively, with a learning rate of 1e-5 and AdamW optimizer [39]. Our camera embedding representation is the Plücker embedding from CameraCtrl, which is computed from the camera’s intrinsic and extrinsic parameters. For real internet data, the intrinsic parameters are set using a heuristic value, while the extrinsic parameters are obtained using SLAM [62] methods. For synthetic data, the intrinsic and extrinsic parameters are directly exported. For the 4 $\times$ batch size setting in stage 1, we use 8 NVIDIA A100 GPUs with gradient accumulation step as 4, or 32 NVIDIA A100 GPUs with gradient accumulation step as 1. The batch size per GPU is always 8 due to the GPU memory limits. We also train the network for 30,000 iterations in 4 $\times$ batch size setting and use the learning rate as 1e-5.
A.2 More Details about Rendering of Synthetic Data
Flow chart of the synthetic data creation pipeline. In Fig. 7, we show a detailed flow chart to help readers better understand our rendering process.
Figure 7: Flow chart of our synthetic data creation pipeline.
Process of anime characters creation. In Fig. 8, we show the illustration of video rendering process in creating anime characters.
A.3 Camera Movement Comparison
As our goal is to train a controllable video diffusion model conditioned on camera trajectory and human pose, we think that let the model be trained on a large range of camera movements (i.e., offsets) in all 6 dimensions (XYZ, roll, yaw, pitch) is important. It is also the major motivation to construct the synthetic part of data in our HumanVid dataset. In Fig. 9, we show statistics of camera pose offsets of these 6 dimensions. Quantitative results shows that the frame-wise camera pose offset (i.e., the camera movement of next frame) of synthetic data part shows larger distribution than the real data part. It illustrates that our camera trajectory design produces large enough camera movement for modeling natural camera movements in the Internet videos. Besides, experimental results on in-the-wild cases in inference shows that model trained on such camera trajectories could follow camera trajectories in real world videos very well.
Besides, the current trajectory of camera is achieved by setting two keyframes: the first frame and the last frame. Although two keyframes could already enable the camera movement to cover a large region, it is also possible to create synthetic videos using three keyframes and five keyframes in our rendering pipeline. Such cases could be found in the supplementary materials or the project website.
Figure 9: Statistics of camera movement strength comparison between synthetic data and real data in the space of location and rotation.
A.4 Visualization of Human Mask, Depth and Normal
Figure 10: Illustration of human segmentation mask (top right), normal map (bottom left), and depth map (bottom right) in our synthetic data.
In Fig. 10, we show the example of paired data of RGB image, human segmentation mask, depth map and normal map. It is achieved by putting SMPL-X characters [10] at a random place of a large-scale city scene dataset [35]. The scene and human are jointly rendered by Unreal Engine. As the major focus of our dataset is not 3D information, we only shows that our rendering pipeline is able to produce human mask, depth and normal, yet our dataset will not provide them.
A.5 Visualization of Human Keypoints Tracking
Figure 11: Illustration of human keypoints tracking by leveraging our synthetic data creation pipeline.
While our rendering pipeline primarily focuses on generating synthetic videos with human pose and camera parameters, it can also perform point tracking functionality similar to PointOdyssey [86]. As demonstrated in Fig. 11, we present a preliminary example of human keypoint tracking, where we specifically track human keypoints rather than all scene or object points. Our rendering pipeline has the potential to support point tracking data generation with some additional development.
A.6 Details of User Study Questionnaires.
We show the description of our user study as follows.
We tested 10 samples, each using 4 different character animation methods.
Please make a comprehensive judgment based on the following factors with
descending importance.
1. Appearance of the character (compared to the first frame) and naturalness
of the movement;
2. Smoothness of the background when the camera is moving;
3. Stability of the image and minimal color flickering.
Please rank the 4 animations for each sample in order of preference.
Then, for each question, we provide the image of first frame, and 4 videos generated by 4 methods. The order of videos are randomized for each question to ensure the fairness of user study.
Appendix B Dataset Documentation
This dataset is designed to train camera-controllable human image animation models. Once trained, these models can also perform previous setting of static camera human image animation by passing camera parameters that represent a static camera to the model. For the latest documentation, please refer to the code in Github [^3] https://github.com/zhenzhiwang/HumanVid.
To train a human image animation model on HumanVid, follow these steps:
  • Download internet videos using the video URLs provided in the text files. Additionally, download the synthetic videos and camera parameters from the cloud drive.
  • Extract human poses from both internet and synthetic videos using DWPose [75].
  • Extract camera trajectories for internet videos using Droid-SLAM [62] following TRAM [69], i.e., using Segment Anything [33] to mask out humans before SLAM.
  • Compile all valid camera parameters, RGB videos, and human pose videos into a JSON file as meta-information.
  • Download pretrained checkpoints for CLIP, SD 1.5, Animatediff, CameraCtrl, etc.
  • Train the first stage of CamAnimate, which involves learning image-level appearance and pose control. The camera encoder at the image level is also trained in this stage.
  • Train the second stage of CamAnimate, which focuses on learning video-level motion representations and camera control capabilities.
Camera Parameter Format. Following Droid-SLAM [62], we use the TUM camera format [^4] https://cvg.cit.tum.de/data/datasets/rgbd-dataset/file_formats to save the camera parameters. Each data line adheres to the following structure: ’timestamp tx ty tz qx qy qz qw’. The timestamp is a floating-point number representing the elapsed seconds since the Unix epoch. The values tx, ty, and tz (float numbers) denote the position of the optical center of the color camera relative to the world origin as defined by the motion capture system. The values qx, qy, qz, and qw (float numbers) represent the orientation of the optical center of the color camera in the form of a unit quaternion, again relative to the world origin as defined by the motion capture system. For camera parameters of internet videos estimated by Droid-SLAM, these are camera-to-world parameters. For synthetic data exported by rendering engines, we set these as world-to-camera parameters. The last two columns of camera parameters for synthetic data represent camera intrinsics, which is the normalized focal length according to the RealEstate10K dataset  [^5] https://google.github.io/realestate10k/download.html, i.e., (focal_length / sensor_width, focal_length / sensor_height).
Appendix C Author Statements
We use the CC-BY 4.0   [^6] https://creativecommons.org/licenses/by/4.0/ license for our HumanVid dataset. We promise that we bear all responsibility in case of violation of rights.
C.1 Licenses
In our dataset, we use assets from the following sources, and we list their licenses below.
BEDLAM [10]: The licensor of Bedlam grants us to use the asset of SMPL and SMPL-X parameters, 3D body and clothing meshes, 2D textures, and scripts for academic usage, according to the statement of their website  [^7] https://bedlam.is.tuebingen.mpg.de/license.html. We use such assets to generate SMPL-X character videos in our HumanVid dataset.
Pexels [5]: Pexels is a copyright-free website that grants us to use their videos for free according to the statement of their website [^8] https://www.pexels.com/license/. We use videos from Pexels as the internet videos part in our dataset.
VroidHub [6]: VroidHub is a platform that collects user-generated contents. The description of license for each 3D asset is in this website [^9] https://hub.vroid.com/en/license?allowed_to_use_user=everyone&characterization_allowed_user=everyone&corporate_commercial_use=allow&credit=necessary&modification=disallow&personal_commercial_use=nonprofit&redistribution=disallow&sexual_expression=allow&version=1&violent_expression=allow. All the 3D avatars we used in our dataset clearly show the permission of usage in their individual websites.
Rokoko [4]: Rokoko is a open-source blender addon to do motion retargeting on animatable 3D avatars. Its license is in this Github Repo  [^10] https://github.com/Rokoko/rokoko-studio-live-blender/blob/master/LICENSE.md, which grants us to use it in our dataset.
Appendix D Maintenance Plan
We will maintain the synthetic videos of our dataset by a cloud drive. For internet videos, we will provide video links to download raw videos from the website, as we could not redistribute these videos.